1. The nth term of an A.P. is given by an = 3 + 4n. The common difference is
(a) 7
(b) 3
(c) 4
(d) 1

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: c Explaination:Reason: We have an = 3 + 4n ∴ an+1 = 3 + 4(n + 1) = 7 + 4n ∴ d = an+1 – an = (7 + 4n) – (3 + 4n) = 7 – 3 = 4